OpenStudy (anonymous):

How do you solve 5y+2=.5(10y+4)

jimthompson5910 (jim_thompson5910):

Start off by distributing 5y+2=0.5(10y+4) 5y+2=0.5(10y)+0.5(4) 5y+2 = 5y + 2 Do you see where to go from here?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

subtract 2 from both sides

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (anonymous):

multiply the right side then bring the values of y to the left side and solve for y!